28 1/2 : 10 1/4 Put this ratio as a fraction in simplest form
Found 2 solutions by stanbon, unlockmath:
Answer by stanbon(75887)

28 1/2 : 10 1/4
----
57/2 / 41/4
-----
= 228/82
-----
= 114: 41
===============
cheers,
Stan H.
===============

Answer by unlockmath(1688)

Hello,
We can make this 28 1/2 : 10 1/4 as:
57/2 / 41/4
Now we can multiply it as:
57/2 * 4/41
228/82 = 114/41 or written as:
2 32/41
Make sense?
RJ
www.math-unlock.com